12 5 R.018 In Hawail, the rate of motor vehicle theft is 792 thefts per 100,000 vehicles. A large parking structure in Honolulu has issued 454 parking permits (a) What is the probability that none of the vehicles with a permit will eventually be stolen? (Round A to 1 decimal place, Use 4 decimal places for your anser (b) What is the probability that at least one of the vehicles with a permit will eventually be stolen? (Use 4 decimal places.) (c) What is the probability that eight or more vehicles with a permit willeventualy be stoten? use & decinal placens) Need Help? Raad h
Answer #1

Ans:

n=484

p=0.00792

Lambda=np=484*0.00792=3.8

a)

P(none)=P(x=0)=e-3.8=0.0224

b)

P(atleast one)=P(x>=1)=1-P(x=0)=1-e-3.8

=1-0.0224=0.9776

c)

P(x>=8)=1-P(x<=7)

=1-POISSON(7,3.8,TRUE)

=0.0401